# Unsourced Adversarial CAPTCHA: A Bi-Phase Adversarial CAPTCHA Framework

### v2

@@ V-A1 Dataset
Imagenet is one of the most commonly used datasets for image classification, image detection, and image localization in deep learning. Although this experiment did not require the images provided by the dataset as the basis for the input, to ensure the experiment’s rigor, we chose to refer to Imagenet with 1000 classifications as input and attacked categories in this experiment.
@@ V-A2 Environment
All experiments are carried out on an Ubuntu 22.04.4 Server with an Intel(R) Xeon(R) Gold 6342 CPU @ 2.80GHz and NVIDIA A40 with 48G of memory.
@@ V-A3 Attack Model
All tasks in this experiment were performed based on eight models, including SeResNeXt101 [25], MobilenetV2 [26], Resnet50 [27], Resnet152, Googlenet [28], Efficientnet [29], inceptionV3 [30] and Alexnet [31].
@@ V-A4 Evaluation Metrics
To demonstrate the effectiveness and superiority of our sparse attack method, we employ the Attack Success Rate (ASR), the Clip Score [32], and the Average Attack Step in comparison with other methods. Specifically, they are defined as:
@@
$\mathrm{ASR}=\frac{N_{adv}}{N_{total}}\times 100\%$ (19)
@@
Clip Score evaluates semantic alignment between images and text using the Clip model, commonly applied in tasks like image generation and text-to-image synthesis.
@@ V-B Comparison and analysis of attack effects
To prove the superior attack performance we demonstrate in white-box scenarios and black-box scenarios, we compare it with classic white-box attack methods such as FGSM [33], BIM [34], and PGD [35], and to prove the superior attack performance of our method, we compare our method with advanced adversarial attack methods based on diffusion model under unknown models, such as U-BigGAN [9], AdvDiffuse [36], DiffAttack [11], AdvDiff [10]. In addition, to prove that BP-UAC can show stable and excellent attack performance against unknown models, we conducted comprehensive experiments on seven traditional DNNs.
As shown in Table IV, we compare the ASR of four classical adversarial attack methods (PGD, FGSM, BIM, and UAC) on seven deep learning models, including the untargeted attack success rate (USR) and the targeted attack success rate (TSR), with the target category “Gila Monster”. The experimental results show that the UAC method exhibits a nearly 100% USR and TSR in all models and scenarios. It significantly outperforms the other attack methods in both USR and TSR, demonstrating its overall robustness and attack effectiveness advantages. In contrast, although the other methods also achieve a high success rate (more than 99%) in untargeted attack scenarios, the ASR is significantly lower in targeted scenarios, especially for the Efficientnet and SeResNext101 models, which suggests that there is still a large room for optimization of the performance of the traditional adversarial attack methods in targeted attacks. The results of this experiment show that untargeted attacks are very mature under the existing methods, and almost all the methods can achieve a very high USR under the white-box untargeted attack scenarios, whereas the effectiveness of targeted attacks is limited by the complexity of the method design and model architecture. The UAC methods perform well in both targeted and untargeted attack scenarios, show good generalization and robustness, and provide a strong benchmark for subsequent research.
@@
As shown in Table V, we have evaluated the ASR using the BP-UAC method on different combinations of baseline and targeted attack models. Overall, the migration ASR of the BP-UAC method is higher than 99% on all model combinations, regardless of the architectural differences between the baseline and target models, which demonstrates the strong generalization ability and robustness of the BP-UAC method in cross-model attacks. Especially on the models of ResNet series, Googlenet, and MobileNetV2, the success rates of the attacks are almost indistinguishable whether they are the baseline model or the target model, which demonstrates that the antiperturbation generated by the BP-UAC method is highly adaptable and stable. It is worth noting that even for models with relatively complex architectures (e.g., SeResNext101 and Efficientnet), the success rate of BP-UAC’s migration attack also remains exceptionally high, further demonstrating its performance advantage in dealing with diverse deep learning models. Overall, the experimental results fully validate the efficiency and robustness of the BP-UAC approach in cross-model scenarios, reflect its potential as a generalized adversarial attack method, and provide critical experimental benchmarks and theoretical support for subsequent research.
As shown in Fig 4, we use the BP-UAC method to compare the migration attack performance with four adversarial attack methods based on generative models using ResNet50 as the baseline model on ResNet152 and InceptionV3 models. From the experimental data, it can be seen that the existing adversarial attack methods based on diffusion model have poor attack performance when facing unknown models, and only the AdvDiff method can barely exceed 50% ASR; in contrast, our method can still show a success rate of close to 100% when facing unknown models, which further illustrates the superiority of our method when facing unknown models. This further demonstrates the superiority of our method in the face of the unknown model. In the production of CAPTCHA, since the illegal model used by the attacker is unknown, therefore, in the production of adversarial CAPTCHA defense, the method is required to have a high degree of robustness in the face of the unknown model. This experiment proves that our method can significantly resist the theft of the CAPTCHA by the illegal attacker.
@@ V-C Ablation study
Figure 6: Comparison plot of adversarial samples generated using detailed prompt versus using labels. It is easy to draw conclusion that: a more detailed prompt will result in a more accurate and higher quality generated image.
We compared adversarial examples generated using class names from ImageNet as prompts with those generated using sentences as prompts, and further incorporated Clip model gradients for optimization. As shown in the first column of Figure 5, although both types of prompts achieved an ASR of 100%, the average number of steps required for a successful attack was noticeably higher when the prompt was a single word. Additionally, the Clip Score of the generated images was lower, indicating poorer quality of the adversarial examples. By incorporating clip model gradients during the generation process, we ensured semantic consistency and visual alignment, which further improved the quality of the generated examples. Overall, the combination of prompt optimization through LLM and gradient optimization via clip model significantly enhanced both attack efficiency and adversarial example quality. Furthermore, as visually demonstrated in Figure 6, when the prompt consisted of a single word, the simplicity and lack of information in the semantic expression led to the generation of incorrect or unrealistic images. In contrast, our optimization strategy effectively guided the generation model to output high-quality images that better aligned with the target class. This comparison clearly demonstrates the crucial role of both LLM and clip model in our method, as they not only improved the quality of the generated examples but also significantly increased the attack efficiency.
@@ V-D Robustness evaluation
In traditional adversarial attack scenarios, various defense preprocessing methods are commonly employed to reduce the efficacy of adversarial samples. However, in practical applications of adversarial CAPTCHA systems, if an attacker preprocesses the CAPTCHA using these defense mechanisms during the attack process, the security of the CAPTCHA is significantly compromised. To validate the robustness of our proposed adversarial CAPTCHA system, this study extensively analyzes the impact of various defense strategies including NRP [37], RS [38], R&P [39] and HGD [40] on adversarial samples. As illustrated in Figure 4, the potency of attack methods based on generative models is typically diminished after defense preprocessing, likely due to increased uncertainty in samples, which reduces the effectiveness of attacks based on generative models. Nonetheless, our BP-UAC, through a bi-path optimization strategy, overcomes the limitations of traditional attack approaches and effectively mitigates the impact of defense measures. By integrating gradient information from multiple models, the BP-UAC enhances the resilience of adversarial CAPTCHAs against a variety of defense mechanisms, while also alleviating the issue of model overfitting.
Figure 7: Adversarial examples generated with different Learning rates.
@@ V-E Influence and setting of Parameters
We also conducted a comprehensive ablation experiment to verify the reasonableness of the parameter settings. We calculated the Clip Score, average attack step and ASR of the adversarial examples generated with different learning rate and different perturb grad scale, as can be seen in the Fig. 7 and the second and third columns of Fig. 5, with the increase of the learning rate, the the average attack step required for the success of the attack decreases significantly, however, a higher learning rate $\eta$ will cause the generative model to be over-guided in some steps during the generation process, thus decreasing the quality of the generated image, for this reason, we consider the quality of the image, the efficiency of the attack, and choose to set $\eta$ to 1.0, and at the same time, for the size of the perturbation $\epsilon$, in order to guarantee the method is effective, we must ensure the success rate of the attack, so in the case of ASR of 100%, we choose to set the perturb grad scale to the $3e^{-3}$ that requires the least number of attack steps and generates the highest Clip Score of the adversarial examples.
@@
In this paper, we propose a novel adversarial CAPTCHA generation framework, BP-UAC, which integrates LLM and generative models while introducing an innovative bi-path adversarial optimization strategy. By overcoming the limitations of traditional adversarial attack methods that add noise to original images, BP-UAC leverages gradients from multiple deep models and simultaneously guides the model toward generating adversarial examples in the directions of the second-highest and target class probability classes. This approach enables BP-UAC to identify more robust paths in different loss spaces, achieving exceptionally high ASR even against unknown models. Experimental results demonstrate that our method not only generates realistic images but also effectively deceives traditional DNN recognition models in various white-box and black-box application scenarios, providing a new direction for future adversarial attack research. Furthermore, given the widespread use of CAPTCHAs in daily life for identity verification and security protection, our method can effectively enhance the security of CAPTCHA systems, preventing unauthorized intrusions and providing essential technical support for CAPTCHA design and security upgrades. In addition, we plan to further explore how to generate higher-quality images in future work.

### v3

@@ V-A1 Dataset
Imagenet is one of the most widely used datasets for image classification, object detection, and object localization in deep learning. Although this experiment did not require the images provided by the dataset as the basis for the input, we refer to Imagenet with 1000 classifications as input and attacked categories in this experiment.
@@ V-A2 Environment
All experiments were conducted on an Ubuntu 22.04.4 Server with an Intel(R) Xeon(R) Gold 6342 CPU @ 2.80 GHz and an NVIDIA A40 with 48 GB of memory.
@@ V-A3 Attack Model
All tasks in this experiment were performed based on eight models, including SeResNeXt101 [41], MobilenetV2 [42], Resnet50 [43], Resnet152, Googlenet [44], Efficientnet [45], inceptionV3 [46], and Alexnet [47].
@@ V-A4 Evaluation Metrics
To demonstrate the effectiveness and superiority of our sparse attack method, we employ the Attack Success Rate (ASR), the Clip Score [48], and the Average Attack Step compared to other methods.
@@
$\mathrm{ASR}=\frac{N_{adv}}{N_{total}}\times 100\%,$ (18)
@@
Clip Score evaluates semantic alignment between images and text using the Clip model, commonly applied in tasks like image generation and text-to-image synthesis:
@@ V-B Comparison and analysis of attack effects
To demonstrate the superior performance of our method in both white-box and black-box scenarios, we conducted comprehensive comparative experiments with DAC and Bi-DAC in both white-box and black-box scenarios.
As shown in Table III, we compare the ASR of ten adversarial attack methods (e.g., SAE [49], Adef [50], ReColorAdv [51], cAdv [52], tAdv [52], ACD [53], ColorFool [54], NCF [55], ACA [56] and DAC) which craft adversarial examples not require the addition of perturbation on seven deep learning models. The experimental results demonstrate that DAC achieves remarkable performance in the black-box adversarial attack task. First, in the face of an unknown model, when the alternative model is different from the target model, the attack success rate of our method far exceeds that of other methods, and is only slightly lower than that of NCF when the alternative model is the same as the target model, which reflects the excellent migratory nature of our method; second, the average attack success rate of our method under two alternative models, RN-50 and MN-v2, reaches 81.9% and 79.75%, which is an improvement of more than 14% over the suboptimal methods, and a significant lead in comprehensive performance; moreover, for complex models, our method shows stronger adaptability, and the advantage is especially prominent in lightweight or high-complexity models, which further proving the superiority of the defensive verification codes generated by our method when facing unknown models.
As shown in Table IV, we compare the ASR of four classical adversarial attack methods (e.g., PGD [57], FGSM [58], BIM [59], and DAC) on seven deep learning models, including the untargeted attack success rate (USR) and the targeted attack success rate (TSR), with the target category “Gila Monster”. The experimental results show that the DAC method exhibits a nearly 100% USR and TSR in all models and scenarios. It significantly outperforms the other attack methods in both USR and TSR, demonstrating its overall robustness and attack effectiveness advantages. In contrast, although the other methods also achieve a high success rate (more than 99%) in untargeted attack scenarios, the ASR is significantly lower in targeted scenarios, especially for the Efficientnet and SeResNext101 models, which suggests that there is still a large room for optimization of the performance of the traditional adversarial attack methods in targeted attacks. Untargeted attacks are highly mature under existing methods, and almost all methods can achieve a very high USR in white-box untargeted attack scenarios. However, the effectiveness of targeted attacks is limited by the complexity of the method design and model architecture. The DAC methods perform well in both targeted and untargeted attack scenarios, show good generalization and robustness, and provide a strong benchmark for subsequent research.
@@
As shown in Table V, we evaluate the ASR using the BP-DAC method on different combinations of baseline and targeted attack models. Overall, the migration ASR of the BP-DAC method is higher than 99% on all model combinations, regardless of the architectural differences between the baseline and target models, which demonstrates the strong generalization ability and robustness of the BP-DAC method in cross-model attacks. On the models of ResNet series, Googlenet, and MobileNetV2, the success rates of the attacks are almost indistinguishable whether they are the baseline model or the target model, which demonstrates that the antiperturbation generated by the BP-DAC method is highly adaptable and stable. It is worth noting that even for models with relatively complex architectures (e.g., SeResNext101 and Efficientnet), the success rate of BP-DAC’s transfer attack remains exceptionally high, further demonstrating its performance advantage in dealing with diverse deep learning models. The experimental results validate the efficiency and robustness of BP-DAC in cross-model scenarios, reflecting its potential as a generalized adversarial attack method and providing critical experimental benchmarks and theoretical support for subsequent research.
As shown in Fig. 4, we use the BP-DAC method to compare the transfer attack performance with four adversarial attack methods based on generative models using ResNet50 as the baseline model on ResNet152 and InceptionV3 models. It can be observed that existing adversarial attack methods based on the diffusion model exhibit poor attack performance against unknown models, and only the AdvDiff method can barely exceed a 50% ASR. In contrast, our method still achieves a success rate of nearly 100% when facing an unknown model. This further demonstrates the superiority of our method in the face of the unknown model. In the production of CAPTCHA, since the illegal model used by the attacker is unknown, the method for adversarial CAPTCHA defense must have a high degree of robustness in the face of the unknown model.
@@ V-C Ablation study
Figure 6: Comparison plot of adversarial samples generated using detailed prompt versus using labels. It is easy to conclude that a more detailed prompt will result in a more accurate and higher-quality generated image.
We compare adversarial examples generated using class names from ImageNet as prompts with those generated using sentences as prompts, and incorporate Clip model gradients for optimization. As shown in the first column of Fig. 5, although both types of prompts achieved an ASR of 100%, the average number of steps required for a successful attack is noticeably higher when the prompt is a single word. Additionally, the Clip Score of the generated images is lower, indicating poorer quality of the adversarial examples. By incorporating clip model gradients during the generation process, we ensure semantic consistency and visual alignment, which further improves the quality of the generated examples. Overall, the combination of prompt optimization through LLM and gradient optimization via the CLIP model significantly enhances both attack efficiency and the quality of adversarial examples. Furthermore, as demonstrated in Fig. 6, when the prompt consists of a single word, the simplicity and lack of information in the semantic expression lead to the generation of incorrect or unrealistic images. In contrast, our optimization strategy effectively guides the generation model to output high-quality images that are better aligned with the target class. This comparison demonstrates the crucial role of both LLM and clip model in our method, as they not only improve the quality of the generated examples but also significantly increase the attack efficiency.
@@ V-D Robustness evaluation
In traditional adversarial attack scenarios, various defense preprocessing methods are commonly employed to reduce the efficacy of adversarial samples. However, in practical applications of adversarial CAPTCHA systems, if an attacker preprocesses the CAPTCHA using these defense mechanisms during the attack process, the security of the CAPTCHA is significantly compromised. To validate the robustness of our proposed adversarial CAPTCHA system, this study extensively analyzes the impact of various defense strategies, including NRP [60], RS [61], R&P [62], and HGD [63], on adversarial samples. As illustrated in Table VI, the potency of attack methods based on generative models is typically diminished after defense preprocessing, likely due to increased uncertainty in the samples, which reduces the effectiveness of attacks based on generative models. Nonetheless, our BP-DAC, through a bi-path optimization strategy, overcomes the limitations of traditional attack approaches and effectively mitigates the impact of defense measures. By integrating gradient information from multiple models, BP-DAC enhances the resilience of adversarial CAPTCHAs against various defense mechanisms, while mitigating the issue of model overfitting.
@@ V-E Influence and setting of Parameters
We also conducted a comprehensive ablation experiment to verify the parameter settings. We calculated the Clip Score, average attack step and ASR of the adversarial examples generated with different learning rate and different perturb grad scale, as show in the second and third columns of Fig. 5, with the increase of the learning rate, the average attack step required for the success of the attack decreases significantly. However, a higher learning rate $\eta$ can cause the generative model to be over-guided in some steps during the generation process, thus decreasing the quality of the generated image. For this reason, we consider the quality of the image, the efficiency of the attack, and choose to set $\eta$ to 1.0. At the same time, for the size of the perturbation $\epsilon$, in order to guarantee the method is effective, we must ensure the success rate of the attack. In the case of ASR of 100%, we choose to set the perturb grad scale to $3\times 10^{-3}$ that requires the least number of attack steps and generates the highest Clip Score of the adversarial examples.
@@
In this paper, we proposed a novel defensive adversarial CAPTCHA generation framework, BP-DAC, which integrates LLM and generative models while introducing an innovative bi-path adversarial optimization strategy. By overcoming the limitations of traditional adversarial attack methods that add noise to original images, BP-DAC leverages gradients from multiple deep models and simultaneously guides the model toward generating adversarial examples in the directions of the second-highest and target class probabilities. This approach enables BP-DAC to identify more robust paths in different loss spaces, achieving exceptionally high ASR even against unknown models. Experimental results demonstrated that our method not only generates realistic images but also effectively deceives traditional DNN recognition models in various white-box and black-box application scenarios, providing a new direction for future adversarial attack research. Furthermore, given the widespread use of CAPTCHAs in daily life for identity verification and security protection, our method can effectively enhance the security of CAPTCHA systems, preventing unauthorized intrusions and providing essential technical support for CAPTCHA design and security upgrades. In the future, we plan to further explore methods for generating higher-quality images.